You currently have $15 and plan to save $5.50 per week. Your friend has $30 and plans to save $3.25 each week. When (at what same time) will you have the same amount of money as your friend?
I have to explain how I solved the problem and show as many stratergies as I can.

Answer by checkley77(12844)

15+5.50X=30+3.25X
5.50X-3.25X=30-15
2.25X=15
X=15/2.25
X=6.667 WEEKS YOU'LL HAVE A MUCH SAVED AS YOUR FRIEND.
PROOF:
15+5.5*6.667=30+3.25*6.667
15+36.67=30+21.67
51.67=51.67